OpenStudy (anonymous):

Patrick deposited $6,875 into a savings account 17 years ago. The account has an interest rate o f 4.9% and the balance is currently $15,734.11. How often does the interest compound? annually, daily, monthly or quarterly??

OpenStudy (anonymous):

fv = pv(1 + i)^N so fv / pv = (1 + i)^N 15734.11 / 6875 = 2.2886 = (1+i)^N If interest was compounded quarterly i = 4.9 / (4 x 100) = 0.01225 and N = 17 x 4 = 68 so (1+i)^N = 1.01225^68 = 2.2886 Since 2.2886 = 2.2886 interest was compounded quartrly. Does this help? :-)